Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 12

1. Section 6.1: problems 5, 55, 63.

2. Section 6.2: problem 39.

3. Section 6.3: problems 21, 25.

4. Section 6.4: problem 5.